# Small time asymptotics of spectral heat content of isotropic processes

### v1

@@ 2.3. Time changed Brownian motion
For any bounded open set $\Omega$ with $C^{1,1}$-boundary we have
$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mathbb{E}[\sqrt{U_{t}}]}=\sqrt{\frac{2}{\pi}}\mathrm{Per}(\Omega).$
@@
Since $U_{t}$ is increasing and continuous, for any $t\geqslant 0$ we have
$\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}B^{(1)}_{U_{t}}\right]=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant U_{t}}B^{(1)}_{s}\right]=\int_{0}^{\infty}\mathbb{E}_{0}\left[\sup_{0\leqslant u\leqslant s}B^{(1)}_{u}\right]\mathbb{P}(U_{t}\in ds).$
Recalling that for any $p>0$,
$\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}B^{(1)}_{s}\right)^{p}\right]=\frac{2^{p/2}t^{p/2}\Gamma(\frac{p+1}{2})}{\sqrt{\pi}}$
we get
(2.4) $\mathbb{E}_{0}\left[\left(\sup_{0\leqslant s\leqslant t}B^{(1)}_{U_{s}}\right)^{p}\right]=\frac{2^{p/2}\Gamma(\frac{p+1}{2})}{\sqrt{\pi}}\mathbb{E}[U^{p/2}_{t}].$
Therefore, (2.3) implies that the conditions of Corollary 1.3 hold. This concludes the proof. ∎
As an example of $U_{t}$, let us consider an $n$-dimensional standard Brownian motion $W=(W_{t})_{t\geqslant 0}$ independent of $B$ and define the integrated squared radial process
$R_{t}=\int_{0}^{t}|W_{s}|^{2}ds.$
Then, the one-dimensional marginal distributions of $X_{t}=B_{R_{t}}$ coincide with the Lévy stochastic area, see [11, Equation (3.5)]. Also, $X$ is neither Markov nor Gaussian. For any $t\geqslant 0$,
$\mathbb{E}[R_{t}]=\int_{0}^{t}nsds=\frac{nt^{2}}{2},$
and by Cauchy-Schwarz inequality $\mathbb{E}[\sqrt{R_{t}}]\leqslant\sqrt{\mathbb{E}[R_{t}]}<\infty$. On the other hand, using Cauchy-Schwarz inequality, we have
$\sqrt{\frac{R_{t}}{t}}\geqslant\frac{1}{t}\int_{0}^{t}|W_{s}|ds,$
@@
$\mathbb{E}[\sqrt{R_{t}}]\geqslant\frac{1}{\sqrt{t}}\mathbb{E}\left[\int_{0}^{t}|W_{s}|ds\right]\geqslant ct$
for some constant $c>0$. Therefore, $\lim_{t\to 0}\mathbb{E}[R_{t}]/\mathbb{E}[\sqrt{R_{t}}]=0$, and hence the conditions in (2.3) are satisfied by $R_{t}$, which implies the following corollary.
Corollary 2.4.
Let the processes $B,R$ be defined as above and $X_{t}=B_{R_{t}}$. Then, for any bounded, open set $\Omega$ with $C^{1,1}$-boundary
$\lim_{t\to 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{\mathbb{E}[\sqrt{R_{t}}]}=\sqrt{\frac{2}{\pi}}\mathrm{Per}(\Omega).$
@@ 3. Preliminaries: Perimeter and regularity of the boundary
for all $x,y\in\mathbb{R}^{d}$. The domain $\Omega\subset\mathbb{R}^{d}$ has $C^{1,\kappa}$-boundary if for every $x\in\partial\Omega$, there exists a neighborhood $V_{x}$ and a $C^{1,\kappa}$ function $\gamma:\mathbb{R}^{d-1}\to\mathbb{R}$ such that
@@
If $\Omega$ has $C^{1}$-boundary, from Stoke’s theorem it follows that $\mathrm{Per}(\Omega)=\mathcal{H}^{d-1}(\partial\Omega)$, the (d1)(d-1)-dimensional Hausdorff measure of $\partial\Omega$.
@@ 4. The minimum functional
In particular, when $f=\mathbbm{1}_{\Omega}$, $Q^{X}_{f}(t)=Q^{X}_{\Omega}(t)$.
@@
as $t\to 0$. Trivially, when $f=\mathbbm{1}_{\Omega}$, $\mathcal{R}^{X}_{f}(t)=\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)$.
@@
$\lim_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx,$
@@
(4.1) $\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\varepsilon\right]=1$
(4.2) $\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\right\}\right]=1$
(4.3) $\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\right\}\right]=1$
(4.4) $\limsup_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\right\}\right]\leqslant 2d$
@@
$\left|\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\varepsilon\right]-\mu(t)\right|\leqslant\max\{\varepsilon,1\}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\wedge 1\right).$
Therefore, (4.1) follows as $\lim_{t\to 0}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\wedge 1\right)/\mu(t)=0$. Next, for any $t>0$, we have
$0$ $\leqslant\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\wedge\varepsilon\right]-\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\right\}\right]$
$\leqslant\varepsilon\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon\right),$
@@
$0$ $\leqslant\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\}\right]-\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]$
$=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon,\sup_{0\leqslant s\leqslant t}|X_{s}|>\varepsilon\}\right]$
$\leqslant\varepsilon\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\varepsilon\right)\leqslant 2d\varepsilon\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}>\varepsilon/\sqrt{d}\right),$
@@
(4.5) $\ \ \ \ \lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]$
$=\lim_{t\to 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\leqslant\varepsilon\}\right]=1.$
@@
$\leqslant 2\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}X^{(i)}_{s}>u\right)=2\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(i)}_{s}\wedge\varepsilon\right].$
@@
$\limsup_{t\to 0}\frac{1}{\mu(t)}\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant 2R\right\}\right]dx\leqslant 2d\varepsilon^{\kappa}\mathrm{Vol}(S_{f}).$
@@
(4.10) $\lim_{t\to 0}\frac{1}{\mu(t)}\left|\mathcal{R}^{X}_{f}(t)-\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{X_{s}\in S_{f},0\leqslant s\leqslant t\}\right]dx\right|=0.$
@@
$=\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}X^{(1)}_{s}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant r\}\right]|\nabla f(x)|=:\mu(t,r)|\nabla f(x)|.$
@@
(4.13) $\limsup_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}$ $\leqslant\limsup_{t\to 0}\frac{\mu(t,2R)}{\mu(t)}\int_{\Omega}|\nabla f(x)|dx$
@@
$\liminf_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}\geqslant\liminf_{t\to 0}\int_{S_{f}}|\nabla f(x)|\frac{\mu(t,\delta(x))}{\mu(t)}dx.$
Since by (4.2) in Lemma 4.3, $\frac{\mu(t,\delta(x))}{\mu(t)}\to 1$ as $t\to 0$, applying Fatou’s lemma we conclude that
(4.14) $\liminf_{t\to 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}\geqslant\int_{S_{f}}|\nabla f(x)|dx=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx.$

### v3

@@ 2.3. Time-changed $\alpha$-stable process
For any bounded open set $\Omega$ with $C^{1,1}$ boundary we have
(2.4) $\lim_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{m(t)}=\mathrm{Per}(\Omega),$
where $m(t)=\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\right$]. In particular when $U_{t}$ has continuous paths, we have $m(t)=\mathbb{E}[\overline{Y}^{(1)}_{1}]\mathbb{E}[U^{1/\alpha}_{t}]$.
Remark 2.4.
In the context of time-changed Lévy processes, SHC of the time-changed process and the killed time-changed process are different in general. This happens as the paths of $U$ are discontinuous in general. However when $U$ has continuous paths, above two notions of SHC coincide since
$\mathbb{P}_{x}\left(\tau^{X}_{\Omega}>t\right)=\mathbb{P}_{x}\left(\tau^{Y}_{\Omega}>U_{t}\right)$
for all $x\in\Omega$ and $t\geqslant 0$. We refer to [9, p. 5] (see also [24, p. 580]) for detailed discussions on the comparison between killed time-changed and time-changed killed processes.
@@
Since $Y$ is an $\alpha$-stable Lévy process, for any $0\leqslant q<\alpha$, $\mathbb{E}[\sup_{0\leqslant s\leqslant t}|Y_{t}|^{q}]<\infty$, see for instance [22, Theorem 25.18]. By $1/\alpha$–self-similarity of $Y$, we have $\mathbb{E}[(\overline{Y}^{(1)}_{t})^{q}]=t^{q/\alpha}\mathbb{E}[(\overline{Y}^{(1)}_{1})^{q}]$. Since $U$ is increasing and right continuous, for any $t\geqslant 0$,
$\overline{X}^{(1)}_{t}\leqslant\sup_{0\leqslant s\leqslant U_{t}}Y^{(1)}_{s}$,
@@
$\mathbb{E}\left[\left(\overline{X}^{(1)}_{t}\right)^{q}\right]\leqslant\mathbb{E}_{0}\left[\left(\overline{Y}^{(1)}_{1}\right)^{q}\right]\mathbb{E}\left[U^{q/\alpha}_{t}\right]<\infty$
for $0\leqslant q<p\wedge\alpha$. On the other hand, using symmetry of the distribution of $X^{(1)}$, we note that for any $u>0$,
$\mathbb{P}_{0}\left(\overline{X}^{(1)}_{t}\geqslant u\right)\geqslant\frac{1}{2}\mathbb{P}_{0}\left(|X^{(1)}_{t}|\geqslant u\right),$
which implies that for any $1\leqslant q<p\wedge\alpha$,
$\mathbb{E}_{0}\left[\left(\overline{X}^{(1)}_{t}\right)^{q}\right]\geqslant\frac{1}{2}\mathbb{E}_{0}\left[|X^{(1)}_{t}|^{q}\right]=\frac{1}{2}\mathbb{E}\left[|Y^{(1)}_{1}|^{q}\right]\mathbb{E}[U^{q/\alpha}_{t}],$
where the last equality follows from self-similarity of $Y$. Therefore, for any $t\geqslant 0$ and $0\leqslant q<p\wedge\alpha$,
$\frac{1}{2}\mathbb{E}_{0}\left[|Y^{(1)}_{1}|^{q}\right]\frac{\mathbb{E}\left[U^{q/\alpha}_{t}\right]}{\mathbb{E}[U^{1/\alpha}_{t}]}\leqslant\frac{\mathbb{E}_{0}\left[\left(\overline{X}^{(1)}_{t}\right)^{q}\right]}{\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\right]}\leqslant\mathbb{E}_{0}\left[\left(\overline{Y}^{(1)}_{1}\right)^{q}\right]\frac{\mathbb{E}\left[U^{q/\alpha}_{t}\right]}{\mathbb{E}[U^{1/\alpha}_{t}]}$
In light of Proposition 1.3, it suffices to show that $\mathbb{E}[U^{q/\alpha}_{t}]/\mathbb{E}[U^{1/\alpha}_{t}]\to 0$ whenever $1<q<p\wedge\alpha$, where $p$ is given by (2.3). By Hölder inequality, for any $1<q<p$ we have
$\frac{\mathbb{E}\left[U^{q/\alpha}_{t}\right]}{\mathbb{E}[U^{1/\alpha}_{t}]}\leqslant\left(\frac{\mathbb{E}\left[U^{p/\alpha}_{t}\right]}{\mathbb{E}[U^{1/\alpha}_{t}]}\right)^{\frac{q-1}{p-1}},$
which shows that the conditions of Proposition 1.3 hold due to (2.3). This concludes the proof of (2.4).
When $U$ has continuous paths, one has
$\sup_{0\leqslant s\leqslant t}Y^{(1)}_{U_{s}}=\sup_{0\leqslant s\leqslant U_{t}}Y^{(1)}_{s}\quad\mbox{$\mathbb{P}_{0}$--almost surely}.$
By independence of $Y$ and $U$, and the self-similarity of $Y$, we conclude that $m(t)=\mathbb{E}[\overline{Y}^{(1)}_{1}]\mathbb{E}[U^{1/\alpha}_{t}]$. ∎
2.3.1. Time-change by inverse subordinator
Consider a subordinator (increasing Lévy process on $(0,\infty)$, see Chapter III in Bertoin [4]) $S=(S_{t})_{t\geqslant 0}$ with $S_{0}=0$ such that
$\mathbb{E}[e^{-\lambda S_{t}}]=e^{-t\varphi(\lambda)}\quad\mbox{for all $\lambda,t>0$},$
where $\varphi$ is a Bernstein function such that
(2.5) $\varphi(\lambda)=b\lambda+\int_{0}^{\infty}(1-e^{-\lambda r})\mu(dr)$
with $b\geqslant 0$, and $\int_{0}^{\infty}\min\{1,r\}\mu(dr)<\infty$. The inverse subordinator $E=(E_{t})_{t\geqslant 0}$ is defined as the right inverse of $(S_{t})_{t\geqslant 0}$, that is,
$E_{t}=\inf\{s\geqslant 0:S_{s}>t\}.$
Then, $E$ is also an increasing process, and it is continuous in time if and only if $S$ has strictly increasing sample paths almost surely, which occurs if and only if $\varphi(\infty)=\infty$.
Proposition 2.5.
Assume that $\varphi$ is regularly varying at infinity with index $\beta\in(0,1)$ and $X_{t}=Y_{E_{t}}$, where $Y$ is a $d$-dimensional isotropic $\alpha$-stable Lévy process with $\alpha\in(1,2]$ and independent of $S$. Then,
$\lim_{t\downarrow 0}\varphi(1/t)^{\frac{1}{\alpha}}(\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t))=\frac{\Gamma(1+1/\alpha)}{\Gamma(1+\beta/\alpha)}\mathbb{E}_{0}\left[\overline{Y}^{(1)}_{1}\right]\mathrm{Per}(\Omega).$
Remark 2.6.
The above proposition has been proved in [9, Theorem 4.4] for $\alpha=2$, and in [8] for $\alpha\in(1,2)$. Below we provide an alternative argument using Proposition 2.3.
Remark 2.7.
When $\varphi(\lambda)=\lambda^{\beta}$, $\beta\in(0,1)$ and $\alpha=2$, $Q^{X}_{\Omega}(t)=\int_{\Omega}u_{\beta}(x,t)dx$, where $u_{\beta}$ is the unique solution to the time-fractional Cauchy problem with Dirichlet boundary condition on $\Omega$, see [16] for details.
Proof.
When $\varphi$ is regularly varying at $\infty$, $\varphi(\infty)=\infty$. Therefore from the discussion above, $E$ has continuous sample paths. By [9, Proposition 4.2], for any $q>0$ one has
(2.6) $\mathbb{E}\left[E^{q}_{t}\right]\sim\frac{\Gamma(q+1)}{\Gamma(q\beta+1)}\varphi(1/t)^{-q}\quad\mbox{as $t\downarrow 0$}.$
Hence, for any $p>1$,
$\frac{\mathbb{E}[E^{p/\alpha}_{t}]}{\mathbb{E}[E^{1/\alpha}_{t}]}\sim\frac{\Gamma(p/\alpha+1)\Gamma(\frac{\beta}{\alpha}+1)}{\Gamma(1+1/\alpha)\Gamma(p\beta/\alpha+1)}\frac{\varphi(1/t)^{1/\alpha}}{\varphi(1/t)^{p/\alpha}}\quad\mbox{as $t\downarrow 0$},$
which shows that $\lim_{t\downarrow 0}\mathbb{E}[E^{p/\alpha}_{t}]/\mathbb{E}[E^{1/\alpha}_{t}]=0$, since $\varphi$ is regularly varying at $\infty$. Therefore using Proposition 2.3 and (2.6) for $q=1/\alpha$, we conclude the proof. ∎
2.3.2. Time-change by any positive, increasing self-similar process
Let $U=(U_{t})_{t\geqslant 0}$ be any positive, $\beta$–self-similar, increasing process with càdlàg paths such that $\mathbb{E}[U^{p/\alpha}_{1}]<\infty$ for some $p>1$, and $Y$ is same as before. Then, $X_{t}=Y_{U_{t}}$ is $\beta/\alpha$–self-similar, that is,
$(cX_{t})_{t\geqslant 0}\overset{d}{=}(X_{tc^{\frac{\alpha}{\beta}}})_{t\geqslant 0}\quad\mbox{for all $c>0$}.$
By self-similarity, we also have
$\mathbb{E}[U^{q/\alpha}_{t}]$ $=t^{q\beta/\alpha}\mathbb{E}[U^{q/\alpha}_{1}]\quad\mbox{for $q\in[0,p]$},$
$\mathbb{E}\left[\overline{X}^{(1)}_{t}\right$] $=t^{\frac{\beta}{\alpha}}\mathbb{E}\left[\overline{X}^{(1)}_{1}\right$].
Therefore, by Proposition 2.3 we deduce the following result.
Proposition 2.8.
For any bounded domain $\Omega$ with $C^{1,1}$ boundary,
$\lim_{t\downarrow 0}\frac{\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)}{t^{\frac{\beta}{\alpha}}}=\mathbb{E}\left[\overline{X}^{(1)}_{1}\right]\mathrm{Per}(\Omega).$
We now provide a specific example of a self-similar increasing process constructed from an increasing self-similar Markov process. Let $\xi=(\xi_{t})_{t\geqslant 0}$ be a positive, increasing, self-similar Markov process with self-similarity index $1/\beta$, that is, for all $c>0$,
$\left(c\xi_{tc^{-\beta}},\mathbb{P}_{x}\right)\overset{d}{=}\left(\xi_{t},\mathbb{P}_{cx}\right)$
By Lamperti’s theorem, see [12], under the probability law $\mathbb{P}_{x}$ for any $x>0$, one can write $\xi$ as a time-changed subordinator as follows:
$\xi_{t}=x\exp\left(S_{A_{tx^{-\beta}}}\right),$
where $S$ is a subordinator and $A_{t}=\inf\{s\geqslant 0:\int_{0}^{s}\exp(\beta S_{r})dr>t\}$. Moreover, the above representation gives a bijection between the class of all positive, increasing, self-similar Markov processes and subordinators. Let $\varphi$ denote the Bernstein function associated with $S$ defined as in (2.5). For simplicity, we exclude the case when $S$ is arithmetic, that is, $b=0$ and $\mu$ is supported on a lattice $r\mathbb{Z}$ for some $r>0$. When $\varphi^{\prime}(0+)<\infty$, or equivalently, $\mathbb{E}[S_{1}]<\infty$, Bertoin and Caballero [5] proved that the process $(\xi_{t},\mathbb{P}_{x})$ converges weakly as $x\downarrow 0$, and therefore, one gets an increasing, self-similar process $\xi$ with $\xi_{0}=0$. Let $\zeta=(\zeta_{t})_{t\geqslant 0}$ denote the right inverse of $\xi$, that is,
$\zeta_{t}=\inf\{s>0:\xi_{s}>t\}.$
Then, $\zeta$ is $\beta$-self-similar, positive, and increasing process but may not be Markovian. From [15, Lemma 2.1] it is known that $\mathbb{E}[\zeta^{p}_{1}]<\infty$ for all $p>0$. For a $d$-dimensional $\alpha$-stable isotropic Lévy process $Y=(Y_{t})_{t\geqslant 0}$ independent of $\xi$, let us define the time-changed process $X_{t}=Y_{\zeta_{t}}$. Then, Proposition 2.8 holds for $X$. We note that $\zeta$ has discontinuous paths whenever $\varphi(\infty)<\infty$.
Since $\varphi^{\prime}(0+)<\infty$, using Fubini theorem and a simple change of variable leads to
$\varphi(\lambda)=b\lambda+\lambda\int_{0}^{1}r^{\lambda-1}m(r)dr,$
where $m$ is a non-decreasing function on $(0,1)$ and $\int_{0}^{1}(-\log r\wedge 1)m(r)dr<\infty$. Patie and Srapionyan [19, Definition 3.1] introduced the following non-local operator operator:
$\partial^{\varphi}_{t}f=t^{1-\beta}bf^{\prime}(t)+t^{-\beta}\int_{0}^{t}f^{\prime}(r)m(r/t)dr.$
In particular, when $\xi$ is the standard $\beta$-stable subordinator with $\beta\in(0,1)$, one has
$b=0\quad\text{and}\quad m(r)=\frac{(1-r)^{-\beta}}{\Gamma(1-\beta)}\mathbbm{1}_{\{0<r<1\}},$
see e.g. [6, Example 5], and in this case $\partial^{\varphi}_{t}$ coincides with the Caputo fractional derivative of order $\beta$. Let us assume that $\zeta$ has continuous paths, or equivalently, $\varphi(\infty)=\infty$. Consider the Cauchy problem
(2.7) $\partial^{\varphi}_{t}u=-(-\Delta_{\Omega})^{\alpha/2}u$
$u(x,0)=\mathbbm{1}_{\Omega}(x),$
where $\Delta_{\Omega}$ is the Dirichlet Laplacian on $\Omega$ and $\alpha\in(1,2]$. Since $(-\Delta_{\Omega})^{\alpha/2}$ is self-adjoint and compact in $L^{2}(\Omega)$, by [19, Theorem 4.1], $u_{\varphi}(x,t)=\mathbb{P}_{x}\left(\tau^{Y}_{\Omega}>\zeta_{t}\right)$ is a solution to the above integro-partial differential equation. Due to path continuity of $\zeta$, we get $u_{\varphi}(x,t)=\mathbb{P}_{x}\left(\tau^{X}_{\Omega}>t\right)$ for all $x\in\Omega$ and $t\geqslant 0$, which shows that the SHC of $X$ is related to the solution to (2.7), that is,
$Q^{X}_{\Omega}(t)=\int_{\Omega}u_{\varphi}(x,t)dx.$
In particular, when $\alpha=2$, $u_{\varphi}$ solves the following Cauchy problem with Dirichlet boundary conditions:
$\partial^{\varphi}_{t}u(x,t)$ $=\Delta_{x}u(x,t)\quad\mbox{in $\Omega\times(0,\infty)$}$
$u(x,t)$ $=0\quad\mbox{for $(x,t)\in\partial\Omega\times(0,\infty)$}$
$u(x,0)$ $=1\quad\mbox{for $x\in\Omega$}.$
@@ 3. Preliminaries: Perimeter and regularity of the boundary
for all $x,y\in\mathbb{R}^{d}$. The domain $\Omega\subset\mathbb{R}^{d}$ has $C^{1,\kappa}$ boundary if for every $x\in\partial\Omega$, there exists a neighborhood $V_{x}$ and a $C^{1,\kappa}$ function $\gamma:\mathbb{R}^{d-1}\to\mathbb{R}$ such that
@@
If $\Omega$ has $C^{1}$ boundary, from Stoke’s theorem it follows that $\mathrm{Per}(\Omega)=\mathcal{H}^{d-1}(\partial\Omega)$, the (d1)(d-1)-dimensional Hausdorff measure of $\partial\Omega$.
@@ 4. The minimum functional
In particular, $Q^{X}_{\mathbbm{1}_{\Omega}}(t)=Q^{X}_{\Omega}(t)$ for any open set $\Omega$.
@@
as $t\downarrow 0$. We note that $\mathcal{R}^{X}_{\mathbbm{1}_{\Omega}}(t)=\mathrm{Vol}(\Omega)-Q^{X}_{\Omega}(t)$.
@@
$\lim_{t\downarrow 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx,$
@@
(4.1) $\lim_{t\downarrow 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\wedge\varepsilon\right]=1$
(4.2) $\lim_{t\downarrow 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\left\{\overline{X}^{(1)}_{t}\leqslant\varepsilon\right\}\right]=1$
(4.3) $\lim_{t\downarrow 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\right\}\right]=1$
(4.4) $\limsup_{t\downarrow 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\sup_{0\leqslant s\leqslant t}|X_{s}|\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\right\}\right]\leqslant 2d$
@@
$\left|\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\wedge\varepsilon\right]-\mu(t)\right|\leqslant\max\{\varepsilon,1\}\mathbb{P}_{0}\left(\overline{X}^{(1)}_{t}>\varepsilon\wedge 1\right).$
Therefore, (4.1) follows as $\lim_{t\downarrow 0}\mathbb{P}_{0}\left(\overline{X}^{(1)}_{t}>\varepsilon\wedge 1\right)/\mu(t)=0$. Next, for any $t>0$, we have
$0$ $\leqslant\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\wedge\varepsilon\right]-\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\left\{\overline{X}^{(1)}_{t}\leqslant\varepsilon\right\}\right]$
$\leqslant\varepsilon\mathbb{P}_{0}\left(\overline{X}^{(1)}_{t}>\varepsilon\right),$
@@
$0$ $\leqslant\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\{\overline{X}^{(1)}_{t}\leqslant\varepsilon\}\right]-\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]$
$=\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\{\overline{X}^{(1)}_{t}\leqslant\varepsilon,\sup_{0\leqslant s\leqslant t}|X_{s}|>\varepsilon\}\right]$
$\leqslant\varepsilon\mathbb{P}_{0}\left(\sup_{0\leqslant s\leqslant t}|X_{s}|>\varepsilon\right)\leqslant 2d\varepsilon\mathbb{P}_{0}\left(\overline{X}^{(1)}_{s}>\varepsilon/\sqrt{d}\right),$
@@
(4.5) $\ \ \ \ \lim_{t\downarrow 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant\varepsilon\}\right]$
$=\lim_{t\downarrow 0}\frac{1}{\mu(t)}\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\{\overline{X}^{(1)}_{t}\leqslant\varepsilon\}\right]=1.$
@@
$\leqslant 2\int_{0}^{\varepsilon}\mathbb{P}_{0}\left(\overline{X}^{(i)}_{t}>u\right)du=2\mathbb{E}_{0}\left[\overline{X}^{(i)}_{t}\wedge\varepsilon\right].$
@@
$\limsup_{t\downarrow 0}\frac{1}{\mu(t)}\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}|X_{s}-x|^{1+\kappa}\mathbbm{1}\left\{\sup_{0\leqslant s\leqslant t}|X_{s}-x|\leqslant 2R\right\}\right]dx\leqslant 2d\varepsilon^{\kappa}\mathrm{Vol}(S_{f}).$
@@
(4.10) $\lim_{t\downarrow 0}\frac{1}{\mu(t)}\left|\mathcal{R}^{X}_{f}(t)-\int_{S_{f}}\mathbb{E}_{x}\left[\sup_{0\leqslant s\leqslant t}-\langle X_{s}-x,\nabla f(x)\rangle\mathbbm{1}\{X_{s}\in S_{f},0\leqslant s\leqslant t\}\right]dx\right|=0.$
@@
$=\mathbb{E}_{0}\left[\overline{X}^{(1)}_{t}\mathbbm{1}\{\sup_{0\leqslant s\leqslant t}|X_{s}|\leqslant r\}\right]|\nabla f(x)|=:\mu(t,r)|\nabla f(x)|.$
@@
(4.13) $\limsup_{t\downarrow 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}$ $\leqslant\limsup_{t\downarrow 0}\frac{\mu(t,2R)}{\mu(t)}\int_{\Omega}|\nabla f(x)|dx$
@@
$\liminf_{t\downarrow 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}\geqslant\liminf_{t\downarrow 0}\int_{S_{f}}|\nabla f(x)|\frac{\mu(t,\delta(x))}{\mu(t)}dx.$
Since by (4.2) in Lemma 4.3, $\frac{\mu(t,\delta(x))}{\mu(t)}\to 1$ as $t\downarrow 0$, applying Fatou’s lemma we conclude that
(4.14) $\liminf_{t\downarrow 0}\frac{\mathcal{R}^{X}_{f}(t)}{\mu(t)}\geqslant\int_{S_{f}}|\nabla f(x)|dx=\int_{\mathbb{R}^{d}}|\nabla f(x)|dx.$